OpenStudy (anonymous):

Solve the nonlinear inequality. Express the solution using interval notation. 4x + 8 over/x − 6 < 0

OpenStudy (anonymous):

\[\frac{4x+8}{x-6}<0\]

OpenStudy (anonymous):

set the numerator = 0 and solve. you get \[4x+8=0\] \[4x=-8\] \[x=-2\]

OpenStudy (anonymous):

set the denominator = 0 and solve. you get \[x-6=0\] \[x=6\]

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

so your fraction will change sign at -2 and at 6

OpenStudy (anonymous):

thats what i was getting. thanks

OpenStudy (anonymous):

from \[(\infty,-2)\] both top and bottom will be negative, and negative over negative is positive

OpenStudy (anonymous):

so what does that look like on the number line

OpenStudy (anonymous):

from \[(-2,6)\] top will be positive but bottom negative and positive over negative is negative.

OpenStudy (anonymous):

and from \[(6,\infty)\] both will be positive. so you now know where this fraction is positive and where it is negative.

OpenStudy (anonymous):

i am not sure what you are asking. what does it look like on a number line. some people draw this

OpenStudy (anonymous):

like brackets or parentheses and which way is line

OpenStudy (anonymous):

4x+8 ----------------(-2)++++++++++++++++++++++++++++++ x-6 ------------------------------------------------(6)+++++++++++++ fraction ++++++++(-2)-------------------------(6) ++++++++++++++

OpenStudy (anonymous):

you want to know where it is negative. it is negative on the interval from -2 to 6

OpenStudy (anonymous):

in interval notation it looks like \[(-2,6)\]

OpenStudy (anonymous):

as a picture on the number line it looks like ______________-2 ################ 6 ______________ where the shaded area is your solution